If a, b, and c are integers and are the sides of a triangle ABC such that b=a(a-c) show that sin(B)/sin(A) is an integer.
Answer by venugopalramana(3286)

If a, b, and c are
integers and are the sides of triangle ABC such that
b=a(a-c), show that sin(B)/sin(A) is an integer
1 solutions
Answer 22049 by venugopalramana(1719)on
2006-05-03 00:49:16 
1.If a, b, and c are integers and are the sides of
triangle ABC such that
> b=a(a-c), show that sin(B)/sin(A) is an integer
IN A TRIANGLE
a/SIN(A)=b/SIN(B)=c/SIN(C)..HENCE
SIN(B)/SIN(A)=b/a...BUT
b=a(a-c)...OR....b/a=a-c...SINCE a AND c ARE INTEGERS
a-c IS AN INTEGER
HENCE b/a=SIN(B)/SIN(A)..IS AN INTEGER.